OpenStudy (anonymous):

write 2 equivalent fractions for 14 over 21

OpenStudy (anonymous):



OpenStudy (anonymous):

14/21=2/3=4/6=8/12=16/24 etc... just do whatever you want to it but make sure you do it to numerator and denominator ex: if you multiply by two multiply both top and bottom by two